OpenStudy (anonymous):

differentiate the following with respect to x: (x^3)(5x + 2)^2 show working please :)

OpenStudy (lalaly):

use the product rule... have u tried to do it?

OpenStudy (anonymous):

yea but the answer given was factorized differently and i didnt understand

OpenStudy (lalaly):

\[\frac{d}{dx}(x^3(5x+2)^2)=3x^2(5x+2)^2+2x^3(5x+2) \times 5\]

OpenStudy (lalaly):

\[=3x^2(5x+2)^2+10x^3(5x+2)\]

OpenStudy (lalaly):

take (5x+2) as a common factor

OpenStudy (lalaly):

\[\huge{(5x+2)(3x^2(5x+2)+10x^3)}\]

OpenStudy (lalaly):

\[\huge{=(5x+2)(15x^3+6x^2+10x^3)}\]

OpenStudy (anonymous):

lalaly is right .

OpenStudy (lalaly):

\[=(5x+2)(15x^3+6x^2)\]

OpenStudy (lalaly):

u can also say\[=x^2(5x+2)(15x+6)\]

OpenStudy (anonymous):

Thank you so much :D

OpenStudy (lalaly):

:D